019> Assessment Stem & Leaf, Histograms, Percentiles Due in 13 hours, 4 Find the 44th percentile, P44、form the following data 103 10.5 11.7 119 12.6 14.9 171 17.7 204 20.6 209 22.4 22.5 23.9 24.3 248 255 25.8 27.3 27.4 29 | 29,5 | 307 | 319 | 347 34.8 36.7 393 43 43.6 43.9 45.2 46.1 46.2 46.7 47.8 48 495 50 sion Hint: Textbook Pages Points possible: 1 This is attempt 1 of 3. Submit howtest.php?
Answer #1

Total number of values = 39

Here we want to find 44th percentile.

The given data-set is already in increasing order.

Next we need to find ( n + 1) *(44/100) = 17.6

The integer part of 17.6 is 17

and the fractional part = 0.6

17th value in the arrange data set = 25.5

18th value = 25.8

Therefore 44th percentile = 25.5 + 0.6*(25.8 - 25.5 ) = 25.5 + (0.6*0.3 ) = 25.5 + 25.5 + 0.18 = 25.68

So final answer is 25.68